OpenStudy (meilendurcer):

-7/9 • 3/5= A. -27/45 B. -7/15 C. 7/15 D. 27/45 i got 20/45.

OpenStudy (anonymous):

in the same manner of the previous exercise .. ( -7 * 3 ) / (9*5)

OpenStudy (anonymous):

Okay, so you have: \[-\frac{ 7 }{ 9 } \times \frac{ 3 }{ 5 }\] What's the first step?

OpenStudy (meilendurcer):

multiply?

OpenStudy (anonymous):

yes.. numerator with numerator and denumerator with denumerator

OpenStudy (meilendurcer):

@Coolsector i did it that way and thats how i got the answer

OpenStudy (anonymous):

ok so -7*3 = -21 and not 20

OpenStudy (anonymous):

I got b. 7/15

OpenStudy (anonymous):

-7/15

OpenStudy (anonymous):

that is right emiko

OpenStudy (meilendurcer):

ahh ok. i did something wrong then.. Thanks